# Primordial magnetogenesis in loop quantum cosmology

### v1

@@ V.4 Backreaction
where $\rho^{\phi}\,=\,\dot{\phi}^{2}/2\,+\,V(\phi)$. The energy densities of magnetic and electric fields are given by Eqns. (6) and (8). They are related to the power spectra through Eqns. (7) and (9). Since the power spectrum of the electric field that is produced with the coupling function Eqn. (15) is much smaller than that of the magnetic field, see figure 1, it is sufficient to check whether $\rho^{\phi}\,>>\,\rho^{B}$. We numerically integrated over the magnetic power spectrum Eqn. (7) to obtain the corresponding energy density at the end of inflation. We find that, at the end of inflation, for the coupling function described by Eqn. (15), and when Minkowski initial conditions are imposed at the bounce
@@
This calculation shows that, for the initial condition and coupling function considered here, the magnetic field can be considered as a test field for the case of the Starobinsky model. The energy density of the magnetic field, though smaller than that of the scalar field, may not be negligible for the case of a quadratic potential. We also computed the ratio of magnetic energy density to energy density of the scalar field for the excited states studied in section V.2. We see that in the case of quadratic potential, the ratio of magnetic energy density to that of scalar field varies from $1.2\,<\,\rho^{B}/\rho^{\phi}\,<\,2.4\times 10^{-7}$, where the higher ratio is obtained when Minkowski initial conditon was imposed at $t\,=\,-5\,{\rm t}_{\rm Pl}$ and lower ratio corresponds to an initial time of $t\,=\,5\,{\rm t}_{\rm Pl}$. For the Starobinsky potential, we find that we obtain significant backreaction when Minkowski initial conditions are imposed before the bounce. When Minkowski initial conditions are imposed after the bounce, say at $t\,=\,10\,{\rm t}_{\rm Pl}$, $\rho^{B}/\rho^{\phi}\,=\,1.7\,\times\,10^{-10}$. Since imposing Minkowski initial conditions at different times is equivalent to imposing excited (non-Minkowski) initial states at the bounce, our calculations show that the energy density of magnetic fields depends on the initial condition. Thus, we see that for both models, there exist initial conditions for which backreaction is negligible.
@@ V.5 Estimate of magnetic field present today
We can substitute the above expression in 28 to obtain $B_{0}$. We find that for the quadratic potential, if Minkowski initial condition is imposed at the bounce, the strength of the primordial magnetic field existing today is $B_{0}\,=\,0.94\,\mu{\rm G}$. For Starobinsky model, value of $B_{0}\,=\,0.08\,\mu{\rm G}$. We also find that if Minkowski initial condition was imposed at $t\,=\,1\,{\rm t}_{\rm Pl}$, $B_{0}\,=3.7\,{\rm nG}$ and $B_{0}\,=\,0.3\,{\rm nG}$ in quadratic and Starobinsky models respectively.
@@
We close this section by considering some cases where scale-dependent features due to the bounce may not be visible today. As explained in section III, the amount of expansion between bounce and the time at which pivot scale leaves the horizon is determined by the value of $\phi_{b}$. We have worked with the minimum value of $\phi_{b}$ which ensures that any departure from scale invariance in primordial scalar power spectrum is only imprinted on wavenumbers smaller than pivot scale, $k_{\star}\,=\,0.002\,{\rm Mpc}^{-1}$. Such a value of $\phi_{b}$ imply $k_{\star}\approx 3\,k_{\rm LQC}$. We see that for this value of $\phi_{b}$, the magnetic power spectrum is also scale independent for modes larger than $k_{\star}$. A larger value of $\phi_{b}$ will blue-shift $k_{\star}$ with respect to $k_{\rm LQC}$. In such a scenario, it is possible that we may not observe the highly scale-dependent part of the magnetic field, and the field we measure will be similar to that generated in slow-roll. We also note that the primordial magnetic field observed today can also become consistent with that generated in slow-roll scenarios for other initial conditions, such as the one corresponding to the Minkowski initial state imposed at $t\,=\,1{\rm t}_{\rm Pl}$.
@@ VI Summary and Discussion
If we consider Minkowski initial states imposed at the time of the bounce, then we find that the magnetic power spectrum, see figure 1, is scale dependent at scales with wavenumbers comparable to or smaller than the scale associated with the bounce. Modes with larger wavenumbers have a nearly scale-invariant magnetic power spectrum. In section V.1, we explained the shape of the power spectrum by analyzing the form of $\sqrt{|f^{\prime\prime}/f|}$ as a function of time. We found that the bounce modifies the state of modes with wavenumbers $k\lesssim k_{b}$ and hence these modes are in an excited state at the onset of inflation. We further verified this by expressing the magnetic power spectrum in terms of Bogoliubov coefficients obtained by relating Minkowski states imposed at a time $t_{\rm mink}$ after the bounce with the states obtained by imposing Minkowski initial conditions at the bounce. We also found that the electric power spectrum generated in this scenario is much smaller than the magnetic power spectrum, as in the case of inflation. We computed the energy density of magnetic fields in section V.4. We find that, in the case of the Starobinsky model, the magnetic energy density contributes only $0.1\%$ to the total energy density. Hence, we conclude that the backreaction is negligible and we can consider the electromagnetic field as a test field in this scenario. On the contrary, in the case of the quadratic potential, we find that magnetic energy density contributes about $14\%$ to the total energy density and hence is not insignificant. In section V.3, we also studied equivalent forms of coupling functions. We find that, in LQC, the coupling functions of the form (21) and (23), which are equivalent to Eqn. (15) during slow-roll, leads to larger magnetic fields. However, the coupling functions written in terms of Ricci scalar Eqn. (24) and scalar field Eqn. (25) are equivalent to Eqn. (15) even in the context of LQC. We would like to reiterate that all the above results were obtained upon imposing Minkowski initial states at the time of bounce.
@@
We also studied the effect of imposing Minkowski initial states at various times around the bounce. Note that this is equivalent to imposing other excited states at the bounce. The magnetic power spectra generated with such initial states for quadratic and Starobinsky potentials are plotted in figures 4 and 5, respectively. We find that the power spectra depend on the time at which initial states are imposed, or equivalently, on the choice of initial states imposed at the bounce. For certain initial states, especially those obtained by imposing Minkowski initial conditions before the bounce, the power spectrum and hence the energy density are larger. This is particularly evident in the case of the Starobinsky model. We find that such initial states can lead to significant backreaction. On the other hand, the power spectra and hence the energy density of modes obtained by imposing the Minkowski initial conditions after the bounce are smaller.
@@
In section V.5, we computed the residual primordial magnetic field observable today. We find that if Minkowski initial states are imposed at the bounce, observable magnetic field, $B_{0}\,=\,0.94\mu{\rm G}$ and $B_{0}\,=\,0.08\mu{\rm G}$ respectively for quadratic and Starobinsky potentials. We also note that if we consider the case where Minkowski initial condition is imposed at t= $1{\rm t}_{\rm Pl}, t$he magnetic field observable today becomes of the order of nano gauss. Planck has constrained the magnetic field observable today to be less than a few nano gauss [80]. We note that the magnetic field obtained upon imposing the Minkowski initial condition at $t\,=\,1{\rm t}_{\rm Pl}$ is consistent with this constraint. Moreover, the observability of the scale-dependent part of the magnetic field depends on the value of $\phi_{b}$. For a larger value of $\phi_{b}$, the features arising due to the bounce will get redshifted and hence may not be observable. In such a case, the magnetic field obtained by imposing the Minkowski initial condition at the bounce may also become viable. Finally, Planck has assumed a power law form of magnetic power spectrum in obtaining this constraint, whereas the scale dependence of the magnetic power spectrum generated in LQC cannot be given by a simple power law. It would be interesting to constrain such a spectrum.

### v2

@@ V.4 Backreaction
where $\rho^{\phi}\,=\,\dot{\phi}^{2}/2\,+\,V(\phi)$. Note that $\rho^{\phi}$ is same as $\rho$ appearing in Eqns. (10). The energy densities of magnetic and electric fields are given by Eqns. (6) and (8). They are related to the power spectra through Eqns. (7) and (9). Since the power spectrum of the electric field that is produced with the coupling function Eqn. (15) is much smaller than that of the magnetic field, see figure 2, it is sufficient to check whether $\rho^{\phi}\,>>\,\rho^{B}$. We numerically integrated over the magnetic power spectrum Eqn. (7) to obtain the corresponding energy density at the end of inflation. We find that, at the end of inflation, for the coupling function described by Eqn. (15), and when Minkowski initial conditions are imposed at the bounce
@@
This calculation shows that, for the initial condition and coupling function considered here, the magnetic field can be considered as a test field for the case of the Starobinsky model. The energy density of the magnetic field, though smaller than that of the scalar field, may not be negligible for the case of a quadratic potential. We also computed the ratio of magnetic energy density to energy density of the scalar field for the excited states studied in section V.2. We see that in the case of quadratic potential, the ratio of magnetic energy density to that of scalar field varies from $1.2\,<\,\rho^{B}/\rho^{\phi}\,<\,2.4\times 10^{-7}$, where the higher ratio is obtained when Minkowski initial condition was imposed at $t\,=\,-5\,{\rm t}_{\rm Pl}$ and lower ratio corresponds to an initial time of $t\,=\,5\,{\rm t}_{\rm Pl}$. For the Starobinsky potential, we find that we obtain significant backreaction when Minkowski initial conditions are imposed before the bounce. When Minkowski initial conditions are imposed after the bounce, say at $t\,=\,10\,{\rm t}_{\rm Pl}$, $\rho^{B}/\rho^{\phi}\,=\,1.7\,\times\,10^{-10}$. Since imposing Minkowski initial conditions at different times is equivalent to imposing excited (non-Minkowski) initial states at the bounce, our calculations show that the energy density of magnetic fields depends on the initial condition. Thus, we see that for both models, there exist initial conditions for which backreaction is negligible.
@@ V.5 Estimate of magnetic field present today
We can substitute the above expression in 28 to obtain $B_{0}$. We find that for the quadratic potential, if Minkowski initial condition is imposed at the bounce, the strength of the primordial magnetic field existing today is $B_{0}\,=\,0.94\,\mu{\rm G}$. For Starobinsky model, value of $B_{0}\,=\,0.08\,\mu{\rm G}$. We also find that if Minkowski initial condition was imposed at $t\,=\,1\,{\rm t}_{\rm Pl}$, $B_{0}\,=3.7\,{\rm nG}$ and $B_{0}\,=\,0.3\,{\rm nG}$ in quadratic and Starobinsky models, respectively.
@@
We close this section by considering some cases where scale-dependent features due to the bounce may not be visible today. As explained in section III, the amount of expansion between bounce and the time at which pivot scale leaves the horizon is determined by the value of $\phi_{b}$. We have worked with the minimum value of $\phi_{b}$ which ensures that any departure from scale invariance in primordial scalar power spectrum is only imprinted on wave numbers smaller than pivot scale, $k_{\star}\,=\,0.002\,{\rm Mpc}^{-1}$. Such a value of $\phi_{b}$ imply $k_{\star}\approx 3\,k_{\rm LQC}$. We see that for this value of $\phi_{b}$, the magnetic power spectrum is also scale independent for modes larger than $k_{\star}$. A larger value of $\phi_{b}$ will blue-shift $k_{\star}$ with respect to $k_{\rm LQC}$. In such a scenario, it is possible that we may not observe the highly scale-dependent part of the magnetic field, and the field we measure will be similar to that generated in slow roll. We also note that the primordial magnetic field observed today can also become consistent with that generated in slow-roll scenarios for other initial conditions, such as the one corresponding to the Minkowski initial state imposed at $t\,=\,1{\rm t}_{\rm Pl}$.
@@ VI Summary and Discussion
If we consider Minkowski initial states imposed at the time of the bounce, then we find that the magnetic power spectrum, see figure 2, is scale dependent at scales with wave numbers comparable to or smaller than the scale associated with the bounce. Modes with larger wave numbers have a nearly scale-invariant magnetic power spectrum. In section V.1, we explained the shape of the power spectrum by analyzing the form of $\sqrt{|f^{\prime\prime}/f|}$ as a function of time. We found that the quantum bounce in LQC changes the shape of $\sqrt{|f^{\prime\prime}/f|}$, which in turn modifies the state of modes with wave numbers $k\lesssim k_{b}$. Hence, these modes are in an excited state at the onset of inflation. These excitations make the power spectrum depart from the nearly scale-invariant behaviour found in slow-roll inflation. We further verified this by expressing the magnetic power spectrum in terms of Bogoliubov coefficients obtained by relating Minkowski states imposed at a time $t_{\rm mink}$ after the bounce with the states obtained by imposing Minkowski initial conditions at the bounce. We also found that the electric power spectrum generated in this scenario is much smaller than the magnetic power spectrum, as in the case of inflation. We computed the energy density of magnetic fields in section V.4. We find that, in the case of the Starobinsky model, the magnetic energy density contributes only $0.1\%$ to the total energy density. Hence, we conclude that the backreaction is negligible and we can consider the electromagnetic field as a test field in this scenario. On the contrary, in the case of the quadratic potential, we find that magnetic energy density contributes about $14\%$ to the total energy density and hence is not insignificant. In section V.3, we also studied equivalent forms of coupling functions. We find that, in LQC, the coupling functions of the form (21) and (23), which are equivalent to Eqn. (15) during slow-roll, leads to larger magnetic fields. However, the coupling functions written in terms of Ricci scalar Eqn. (24) and scalar field Eqn. (25) are equivalent to Eqn. (15) even in the context of LQC. We would like to reiterate that all the above results were obtained upon imposing Minkowski initial states at the time of bounce.
@@
We also studied the effect of imposing Minkowski initial states at various times around the bounce. Note that this is equivalent to imposing other excited states at the bounce. The magnetic power spectra generated with such initial states for quadratic and Starobinsky potentials are plotted in figures 5 and 6, respectively. We find that the power spectra depend on the time at which initial states are imposed, or equivalently, on the choice of initial states imposed at the bounce. For certain initial states, especially those obtained by imposing Minkowski initial conditions before the bounce, the power spectrum and hence the energy density are larger. This is particularly evident in the case of the Starobinsky model. We find that such initial states can lead to significant backreaction. On the other hand, the power spectra and hence the energy density of modes obtained by imposing the Minkowski initial conditions after the bounce are smaller.
@@
Finally, in section V.5, we computed the residual primordial magnetic field observable today. We find that if Minkowski initial states are imposed at the bounce, observable magnetic field, $B_{0}\,=\,0.94\mu{\rm G}$ and $B_{0}\,=\,0.08\mu{\rm G}$ respectively for quadratic and Starobinsky potentials. We also note that if we consider the case where Minkowski initial condition is imposed at t= $1{\rm t}_{\rm Pl}, t$he magnetic field observable today becomes of the order of nano gauss. Planck has constrained the magnetic field observable today to be less than a few nano gauss [85]. We note that the magnetic field obtained upon imposing the Minkowski initial condition at $t\,=\,1{\rm t}_{\rm Pl}$ is consistent with this constraint. Moreover, we also noted that the observability of the scale-dependent part of the magnetic field depends on the value of $\phi_{b}$. For a larger value of $\phi_{b}$, the features arising due to the bounce will get redshifted and hence may not be observable. In such a case, the magnetic field obtained by imposing the Minkowski initial condition at the bounce may also become viable.
As mentioned in the introduction, there are at least two reasons for studying magnetogenesis in the context of LQC. To understand if any or what signatures of the quantum bounce will be carried by the magnetic fields, and to find out whether magnetic fields generated in LQC are consistent with the current constraints on primordial magnetic fields. Our analysis shows that the primary effect of the quantum bounce in LQC is to introduce a strong scale dependence in the primordial magnetic field as opposed to the nearly scale-invariant magnetic field generated in slow-roll inflation. The quantum bounce modifies the initial states of the modes before they leave the horizon during inflation. These excitations manifest as scale dependence in the magnetic power spectrum. Our analysis also reveals that, despite the potential, there are initial states that can lead to scale-dependent magnetic fields, which are in agreement with the observable constraints. This work, thus, opens an intriguing possibility, namely, the ability to investigate quantum bounce in LQC using primordial magnetic fields. There are several avenues one could explore.
For instance, there have been proposals for initial conditions based on fundamental principles for primordial scalar perturbations at or around the bounce [86, 87, 88, 89, 90]. It would be interesting to extend these proposals to the vector potential and investigate the nature of the magnetic field generated with these initial conditions in LQC. Such investigations can help us arrive at the principles that could lead us to a well-motivated initial state. Second, it is known that anisotropies in the background spacetime grow as the universe contracts. Hence, it is possible that spacetime at the time of the bounce is anisotropic. Since the inflationary expansion that follows the bounce dilutes this anisotropy, today’s background spacetime will be isotropic. However, the primordial perturbations retain the imprints of this anisotropic phase [69, 70]. It would be interesting to explore the effects of anisotropic quantum bounce on primordial magnetic fields. Moreover, the presence of anisotropy during the quantum regime breaks the conformal invariance of the electromagnetic action. This may provide an alternate mechanism for magnetogenesis in LQC. Moreover, Planck had assumed a power-law form of the magnetic power spectrum in obtaining constraints on magnetic fields, whereas the scale dependence of the magnetic power spectrum generated in LQC cannot be described by a simple power law. Analysing imprints of such a spectrum could help us in arriving at more realistic constraints.
Acknowledgement
We thank Ivan Agullo and L. Sriramkumar for their comments.